# A comprehensive study of the relations between the properties of planetary systems and the chemical compositions of their host stars

### v1

@@ Abstract
The giant planet–metallicity correlation revealed that planetary formation depends on the stellar properties. There is growing evidence that it is also valid for smaller hot planets, but it is not clear whether elements other than iron also influence the properties of planetary systems. To investigate this, we determined the abundances of 13 chemical elements (Na, Mg, Al, Si, Ca, Sc, Ti, V, Cr, Mn, Co, Ni and Cu) for a sample of 561 Kepler exoplanet-hosting stars using high-resolution Keck/HIRES spectra. We find that stars in systems having only large or hot planets are enriched in some elements relative to those having only small or warm planets, respectively, with this signature being related to the underlying stellar metallicity. This Kepler sample is composed of stars belonging to the Galactic low- and high- sequences, corresponding to the chemical thin and thick disks. Our results reveal that stars enhanced in -elements may facilitate the formation of large planets in metal-poor environments although the iron abundance is still a limiting factor. We also investigated chemical abundances as a function of elemental condensation temperatures and found that there is a diversity of slopes of regardless of the exoplanetary systems hosted by the star. We confirmed that the Sun is depleted in refractory elements relative to the solar twins in our sample, all of which host a diversity of exoplanets, suggesting that this depletion is caused by processes not related to planet formation.
@@ I Introduction
The formation, evolution, architectures and habitability of planetary systems depend on the physical properties and chemical compositions of their host stars. The strongest evidence for this connection is the planet-metallicity correlation, first proposed by gonzalez97 and subsequently confirmed by multiple studies (e.g., santos04; fischer05; ghezzi10; ghezzi18). Independent analyzes of different samples consistently showed that the frequency of giant planets increases with the stellar metallicity for FGK main-sequence stars. This relation provides strong support for the core accretion mechanism of planet formation (il04).
@@
Follow-up studies investigated whether this correlation is also valid for smaller planets. Due to technical limitations at the time, these were based on small samples but revealed a different picture. Stars hosting only Neptunian-mass planets were not preferentially metal-rich (e.g., sousa08; ghezzi10). Leveraging the much larger sample of stars with small planets discovered by the Kepler mission (borucki10), buchhave12 showed that planets with orbit stars with a wide range of metallicities, while giant planets are preferentially found around more metal-rich hosts (but see also wf15). In a subsequent study, buchhave14 classified 600 exoplanet candidates into three different groups according to the metallicity distributions of their 400 parent stars, suggesting that it is an important parameter in shaping the architectures of planetary systems.
@@
Using a sample of more than 20,000 Kepler stars that included 665 planet candidates, mulders16 further refined this result showing that the occurrence rates of hot exoplanets ( 10 days) were almost three times higher for stars that are more metal-rich than the Sun compared to their more metal-poor counterparts. wilson18 found a similar separation from the analysis of 282 Kepler Objects of Interest (KOIs) observed by the Sloan Digital Sky Survey (SDSS) IV Apache Point Observatory Galactic Evolution Experiment (APOGEE; majewski17), although for a slightly smaller orbital period ( 8 days).
@@
Analyzing an even larger sample of 1305 planet-hosting stars from the California-Kepler Survey (CKS; petigura17), petigura18 concluded that the occurrence rate of exoplanets depends on stellar metallicity, but with varying strengths according to the planetary radii and orbital periods. For instance, the occurrence of super-Earths () increases with stellar metallicities for the hot planets ( 10 days) but not for their warm counterparts (10 days 100 days). Also using a sample taken from the CKS, ghezzi21 found a statistically significant difference between the metallicity distributions of systems with single hot super-Earths and single warm super-Earths. However, a similar result was not observed for systems with multiple super-Earths or with single or multiple sub-Neptunes (). These efforts revealed that the planet-metallicity correlation is much more complex than initially thought (see also adibekyan19; teske24).
@@
Complementary studies investigated whether elements other than iron might also be related to the process of planetary formation. For instance, robinson06 found that stars hosting giant planets were more enriched in silicon and nickel relative to a population of metal-rich stars without known planets. brugamyer11 provided additional support for the dependence of the occurrence rate of giant planets on stellar silicon abundances, however they did not find a similar result for oxygen. adibekyan12a; adibekyan12b presented evidence that planet occurrence is higher among stars with larger abundances of -elements in the metal-poor regime, suggesting that other elements are important for planetary formation when small amounts of iron are available.
@@
More recently, wilson22 analyzed a sample of 1018 KOIs that also had chemical abundances for ten elements (C, Mg, Al, Si, S, K, Ca, Mn, Fe, and Ni) from APOGEE (majewski17). They found that the occurrence rates of hot super-Earths () and sub-Neptunes () increase for higher abundances of any of the elements, with a stronger correlation for the latter class. These results extended those of petigura18 for nine additional elements.
@@
Instead of focusing on individual elements, torres-quijano25 applied a machine-learning algorithm to an updated version of the Hypatia Catalog (hinkel14) to search for abundance patterns that could indicate the presence of small planets (). Surprisingly, different experiments revealed that sodium (Na) and vanadium (V) were always among the most important elements and that aluminum (Al) also seemed to be relevant (in agreement with the results of wilson22, for Al).
@@
Current evidence from multiple studies thus suggests that the chemical abundances of different elements influence the formation of planets. However, this process might also affect the chemical composition of the host stars. melendez09 showed that the Sun is depleted in refractory elements (i.e., with condensation temperatures 900 K) relative to the solar twins. They suggested that this could be a signature of the formation of rocky planets, since the “missing” material would be locked up in the terrestrial planets. schuler15 investigated this hypothesis using a sample of seven stars with at least one small planet () and did not find the signature for any of them. Subsequent studies based on larger samples revealed that although the Sun is indeed depleted in refractory elements compared to solar analogs or twins, this difference does not seem to be related to planetary formation, since stars with planets exhibit a diversity of trends of elemental abundances as a function of condensation temperature (e.g., bedell18; liu20; nibauer21; rampalli24; carlos25; sun25a; sun25b; rampalli25; see also the earlier study from smith01 and the recent review by gustafsson25). However, the samples of planet-hosting stars analyzed were relatively small (50), while the sample of stellar hosts analyzed in this study is much larger and can offer further insights into this question.
@@
In this paper, we build upon these previous works and that of ghezzi21 to further investigate possible correlations between the formation and architectures of planetary systems and detailed chemical abundances of their host stars. We use a subsample from the CKS since its large size ( 500 planet-hosting stars) allows statistical conclusions to be drawn for a variety of planetary systems. We provide independent determinations of the abundances of 13 elements (Na, Mg, Al, Si, Ca, Sc, Ti, V, Cr, Mn, Co, Ni, Cu), three of which (Sc, Co and Cu) were never analyzed before for this sample. This paper is organized as follows. In Section II, we describe our sample, the data used, and the determination of chemical abundances and stellar evolutionary parameters. We present our results, perform consistency checks, and compare them with the literature in Section III. We discuss our findings in Section IV and present our concluding remarks in Section V.
@@ II.1 Sample and Data
We selected our sample from ghezzi21, which originally consisted of 663 planet-hosting stars. This was the “clean” sample obtained by martinez19 from the CKS catalog (petigura17). However, ghezzi21 removed 102 stars that had large () median uncertainties on the equivalent widths measured for the iron lines and/or microturbulence velocities greater than 1.7 km s-1. These quality cuts provided a final sample of 561 planet-hosting stars that will be analyzed in this study. Their distances, retrieved from bj21, range from 60 to 1800 pc, with a median value of 630 pc. For this sample, we adopt the precise and homogeneous stellar parameters derived in martinez19 and ghezzi21, shown in Table 1 and the Kiel diagram presented in Figure 1.
@@
The planet radii for the sample were calculated by martinez19 and the corresponding planetary systems were classified by ghezzi21 according to all confirmed planets orbiting the host star. Briefly, planets with , , and are considered super-Earths (SE), sub-Neptunes (SN), sub-Saturns (SS) and Jupiters (JP), respectively. Planets with orbital periods 10 days and 10 days days are classified as hot (H) and warm (W), respectively. Finally, systems with only one or more than one detected planet are considered, respectively, single or multiple. For instance, a system classified as “SE single H” has one hot super-Earth while one in the category “SN multi W” has at least two warm sub-Neptunes.
The spectra are the same as analyzed by ghezzi21 and were obtained by the CKS team (petigura17) with the High Resolution Echelle Spectrometer (HIRES; vogt94) spectrograph attached to the Keck I 10 m telescope (Maunakea, Hawaii). The spectra have high resolution (R 60,000), almost complete wavelength coverage between 3640 Å and 7990 Å and are publicly available111https://california-planet-search.github.io/cks-website/. Their signal-to-noise (S/N) ratios vary from 20 to 300 and have a median value of 60, as measured by ghezzi21.
@@ II.2 Line List
We compiled an initial list of 231 lines for the 16 species analyzed in this work (Na I, Mg I, Al I, Si I, Ca I, Sc I, Sc II, Ti I, Ti II, V, Cr I, Cr II, Mn I, Co I, Ni I, Cu I) from the following sources: melendez14, bedell14, schuler15, teske15, bedell18 and liu18. Their atomic parameters were retrieved from the Vienna Atomic Line Database (piskunov95; ryabchikova97; kupka99; kupka00; ryabchikova15; pakhomov19). For Sc, V, Mn, Co and Cu, we considered the hyperfine structure and isotopic splittings. We visually inspected all lines in the Solar Atlas spectrum from hinkle00, which is the one originally published by kurucz84 but corrected for telluric absorption (melendez06), and removed eight lines that were too weak, blended or close to strong lines (e.g., H).
We automatically measured the equivalent widths (EWs) for the remaining lines using ARES v2 (sousa07; sousa15) with the following input parameters: smoothder = 4, space = 3.0, rejt = 0.999, lineresol = 0.1 and miniline = 5. We noted that ARES did not provide good fits for some lines, so we manually measured their EWs using the splot task in the Community Distribution of IRAF (tody86; tody93; iraf_ascl; iraf_zenodo)222Image Reduction and Analysis Facility (IRAF) was written by the National Optical Astronomy Observatories (NOAO) and is now maintained by the iraf-community. IRAF is listed in the Astronomical Source Code Library as ascl:9911.002. The DOI is 10.5281/zenodo.5816743..
Using this revised line list and the 2019 version of the software MOOG (sneden73)333https://www.as.utexas.edu/~chris/moog.html, we determined the chemical abundances for the Sun. The model atmosphere was adopted from the KURUCZ ATLAS9 ODFNEW grid (ck03) and has the following parameters: = 5777 K, = 4.44, [Fe/H] = 0.00 and = 1.00 km s-1. We used the driver blends on MOOG for Sc, V, Mn, Co and Cu and abfind for the other elements, both with option 1 for the treatment of damping, i.e., van der Waals damping factors within the code (barklem00; barklem05) are adopted if available, otherwise values from our line list are utilized. After checking the results, we adjusted the gf values so that all lines for a given element returned the corresponding solar abundances from asplund09. Finally, we removed 18 lines that are located in the inter-order or inter-CCD spaces in the HIRES spectra. Our final list contains 205 lines and can be seen in Table 2, which also contains the measured EWs for the Solar Atlas.
@@ II.3 Chemical Abundances
The chemical abundances for the stars in our sample were obtained in a similar manner as for the Solar Atlas. The model atmospheres were interpolated from the KURUCZ ATLAS9 ODFNEW grid (ck03) using the atmospheric parameters (, [Fe/H] and ) from martinez19 and ghezzi21. The EWs were automatically measured with ARES v2 (sousa07; sousa15) with the same input parameters described in Section II.2, except for the rejt parameter, which was replaced by the S/N value of the spectrum of each star.
@@
Note, however, that EWs could not be measured for all lines because some of them were too weak or blended, depending on the stellar parameters and S/N of the spectra. In particular, ARES was not able to measure any EWs for Sc I for seven stars: KOI-244, KOI-262, KOI-1530, KOI-2833, KOI-3165, KOI-3438 and KOI-3928. They are hotter stars with lower metallicities, so Sc I lines are very weak (EW 5 mÅ).
@@
For each star, we removed lines with measured depths lower than zero or higher than one, as these are non-physical values. Additionally, we cut lines with relative errors on the EWs larger than 50%, which indicate poor fits to the profiles, or with EW 150 mÅ, to avoid very strong lines. The cut based on the relative errors left no Al lines for the stars KOI-49, KOI-620, KOI-1779 and KOI-2022. All EWs used for the following abundance determination are provided in Table 2.
@@
We determined the chemical abundances using the driver blends of the 2019 version of MOOG (sneden73) for elements that have hyperfine structure and isotopic splittings (Sc, V, Mn, Co and Cu) and abfind otherwise, both with option 1 for the damping input parameter. For species that had two or more lines, we made an initial cut of lines having abundances with absolute deviations from the median larger than 0.40 dex to remove lines with possible problems in their EW measurements. This procedure left no Na lines for the star KOI-837. Then, we performed a cut of lines for which abundances were outside a 2-MAD (median absolute deviation) interval around the median. However, we noted that this procedure was not enough to remove bad measurements in all cases, so we added two additional clippings to remove abundances with absolute deviations greater than 0.20 dex. This iterative process ensured that we ended up with precise abundances even for stars with the lowest S/N spectra. Finally, we subtracted the mean abundances from the reference solar values given by asplund09 to obtain [X/H] values, where X is a given element. The abundances for all stars in our sample are shown in Table 1, which also includes the values obtained from the analysis of the spectrum of sunlight reflected off Vesta (see Section III.2.2).
@@
We calculated the abundance uncertainties considering the contributions from the dispersions in the abundances returned by the individual lines as well as the variations caused by the uncertainties in each of the four atmospheric parameters (, , [Fe/H] and ). The former was estimated from the standard deviation of the mean for the absolute abundance A(X), if the element X had abundance measurements from two or more lines. The contribution of the effective temperature for the uncertainty was obtained by generating two new model atmospheres, with (), and determining two new sets of abundances for all elements. Then, we calculated the differences A(X)final - A(X) and the maximum absolute value was taken as the uncertainty caused by the error on . A similar procedure was done for estimating the uncertainties caused by the errors in , [Fe/H] and . The total uncertainties for the abundances of each element were determined by adding all five contributions in quadrature and are shown in Table 1.

### v2

@@ Abstract
The giant planet–metallicity correlation revealed that planetary formation depends on the stellar properties. There is growing evidence that it is also valid for smaller hot planets, but it is not clear whether elements other than iron also influence the properties of planetary systems. To investigate this, we determined the abundances of 13 chemical elements (Na, Mg, Al, Si, Ca, Sc, Ti, V, Cr, Mn, Co, Ni, and Cu) for a sample of 561 Kepler exoplanet-hosting stars using high-resolution Keck/HIRES spectra. We find that stars in systems having only large or hot planets are enriched in some elements relative to those having only small or warm planets, respectively, with this signature being related to the underlying stellar metallicity. This Kepler sample is composed of stars belonging to the Galactic low- and high- sequences, corresponding to the chemical thin and thick disks. Our results reveal that stars enhanced in -elements may facilitate the formation of large planets in metal-poor environments although the iron abundance is still a limiting factor. We also investigated chemical abundances as a function of elemental condensation temperatures and found that there is a diversity of slopes regardless of the exoplanetary systems hosted by the star. We confirmed that the Sun is depleted in refractory elements relative to the solar twins in our sample, all of which host a diversity of exoplanets, suggesting that this depletion is caused by processes not related to planet formation.
@@ I Introduction
The formation, evolution, architectures, and habitability of planetary systems depend on the physical properties and chemical compositions of their host stars. The strongest evidence for this connection is the planet-metallicity correlation, first proposed by Gonzalez (1997) and subsequently confirmed by multiple studies (e.g., Santos et al., 2004; Fischer and Valenti, 2005; Ghezzi et al., 2010, 2018). Independent analyses of different samples consistently showed that the frequency of giant planets increases with the stellar metallicity for FGK main-sequence stars. This relation provides strong support for the core accretion mechanism of planet formation (Ida and Lin, 2004).
@@
Follow-up studies investigated whether this correlation is also valid for smaller planets. Due to technical limitations at the time, these were based on small samples but revealed a different picture. Stars hosting only Neptunian-mass planets were not preferentially metal rich (e.g., Sousa et al., 2008; Ghezzi et al., 2010). Leveraging the much larger sample of stars with small planets discovered by the Kepler mission (Borucki et al., 2010), Buchhave et al. (2012) showed that planets with orbit stars with a wide range of metallicities, while giant planets are preferentially found around more metal-rich hosts (but see also Wang and Fischer, 2015). In a subsequent study, Buchhave et al. (2014) classified 600 exoplanet candidates into three different groups according to the metallicity distributions of their 400 parent stars, suggesting that it is an important parameter in shaping the architectures of planetary systems.
@@
Using a sample of more than 20,000 Kepler stars that included 665 planet candidates, Mulders et al. (2016) further refined this result showing that the occurrence rates of hot exoplanets ( 10 days) were almost 3 times higher for stars that are more metal rich than the Sun compared to their more metal-poor counterparts. Wilson et al. (2018) found a similar separation from the analysis of 282 Kepler Objects of Interest (KOIs) observed by the Sloan Digital Sky Survey (SDSS) IV Apache Point Observatory Galactic Evolution Experiment (APOGEE; Majewski et al., 2017), although for a slightly smaller orbital period ( 8 days).
@@
Analyzing an even larger sample of 1305 planet-hosting stars from the California-Kepler Survey (CKS; Petigura et al., 2017), Petigura et al. (2018) concluded that the occurrence rate of exoplanets depends on stellar metallicity, but with varying strengths according to the planetary radii and orbital periods. For instance, the occurrence of super-Earths () increases with stellar metallicities for the hot planets ( 10 days) but not for their warm counterparts (10 days 100 days). Also using a sample taken from the CKS, Ghezzi et al. (2021) found a statistically significant difference between the metallicity distributions of systems with single hot super-Earths and single warm super-Earths. However, a similar result was not observed for systems with multiple super-Earths or with single or multiple sub-Neptunes (). These efforts revealed that the planet-metallicity correlation is much more complex than initially thought (see also Adibekyan, 2019; Teske, 2024).
@@
Complementary studies investigated whether elements other than iron might also be related to the process of planetary formation. For instance, Robinson et al. (2006) found that stars hosting giant planets were more enriched in silicon and nickel relative to a population of metal-rich stars without known planets. Brugamyer et al. (2011) provided additional support for the dependence of the occurrence rate of giant planets on stellar silicon abundances, however they did not find a similar result for oxygen. Adibekyan et al. (2012b, a) presented evidence that planet occurrence is higher among stars with larger abundances of -elements in the metal-poor regime, suggesting that other elements are important for planetary formation when small amounts of iron are available.
@@
More recently, Wilson et al. (2022) analyzed a sample of 1018 KOIs that also had chemical abundances for 10 elements (C, Mg, Al, Si, S, K, Ca, Mn, Fe, and Ni) from APOGEE (Majewski et al., 2017). They found that the occurrence rates of hot super-Earths () and sub-Neptunes () increase for higher abundances of any of the elements, with a stronger correlation for the latter class. These results extended those of Petigura et al. (2018) for nine additional elements.
@@
Instead of focusing on individual elements, Torres-Quijano et al. (2025) applied a machine learning algorithm to an updated version of the Hypatia Catalog (Hinkel et al., 2014) to search for abundance patterns that could indicate the presence of small planets (). Surprisingly, different experiments revealed that sodium (Na) and vanadium (V) were always among the most important elements and that aluminum (Al) also seemed to be relevant (in agreement with the results of Wilson et al., 2022, for Al).
@@
Current evidence from multiple studies thus suggests that the chemical abundances of different elements influence the formation of planets. However, this process might also affect the chemical composition of the host stars. Meléndez et al. (2009) showed that the Sun is depleted in refractory elements (i.e., with condensation temperatures 900 K) relative to the solar twins. They suggested that this could be a signature of the formation of rocky planets, since the “missing” material would be locked up in the terrestrial planets. Schuler et al. (2015) investigated this hypothesis using a sample of seven stars with at least one small planet () and did not find the signature for any of them. Subsequent studies based on larger samples revealed that, although the Sun is indeed depleted in refractory elements compared to solar analogs or twins, this difference does not seem to be related to planetary formation, since stars with planets exhibit a diversity of trends of elemental abundances as a function of condensation temperature (e.g., Bedell et al. 2018; Liu et al. 2020; Nibauer et al. 2021; Rampalli et al. 2024; Martos et al. 2025; Carlos et al. 2025; Sun et al. 2025b, a; Rampalli et al. 2025; see also the earlier study from Smith et al. 2001 and the recent review by Gustafsson 2025). However, the samples of planet-hosting stars analyzed were relatively small (50), while the sample of stellar hosts analyzed in this study is much larger and can offer further insights into this question.
@@
In this paper, we build upon these previous works and that of Ghezzi et al. (2021) to further investigate possible correlations between the formation and architectures of planetary systems and detailed chemical abundances of their host stars. We use a subsample from the CKS since its large size ( 500 planet-hosting stars) allows statistical conclusions to be drawn for a variety of planetary systems. We provide independent determinations of the abundances of 13 elements (Na, Mg, Al, Si, Ca, Sc, Ti, V, Cr, Mn, Co, Ni, Cu), three of which (Sc, Co, and Cu) were never analyzed before for this sample. This paper is organized as follows. In Section II, we describe our sample, the data used, and the determination of chemical abundances and stellar evolutionary parameters. We present our results, perform consistency checks, and compare them with the literature in Section III. We discuss our findings in Section IV and present our concluding remarks in Section V.
@@ II.1 Sample and Data
We selected our sample from Ghezzi et al. (2021), which originally consisted of 663 planet-hosting stars. This was the “clean” sample obtained by Martinez et al. (2019) from the CKS catalog (Petigura et al., 2017). However, Ghezzi et al. (2021) removed 102 stars that had large () median uncertainties on the equivalent widths (EWs) measured for the iron lines and/or microturbulence velocities greater than 1.7 km s-1. These quality cuts provided a final sample of 561 planet-hosting stars that will be analyzed in this study. Their distances, retrieved from Bailer-Jones et al. (2021), range from 60 to 1800 pc, with a median value of 630 pc. For this sample, we adopt the precise and homogeneous stellar parameters derived in Martinez et al. (2019) and Ghezzi et al. (2021), shown in Table 1 and the Kiel diagram presented in Figure 1.
@@
The planet radii for the sample were calculated by Martinez et al. (2019), and the corresponding planetary systems were classified by Ghezzi et al. (2021) according to all confirmed planets orbiting the host star. Briefly, planets with , , and are considered super-Earths (SE), sub-Neptunes (SN), sub-Saturns (SS) and Jupiters (JP), respectively. Planets with orbital periods 10 days and 10 days days are classified as hot (H) and warm (W), respectively. Finally, systems with only one or more than one detected planet are considered, respectively, single or multiple. For instance, a system classified as “SE single H” has one hot super-Earth while one in the category “SN multi W” has at least two warm sub-Neptunes.
The spectra are the same as analyzed by Ghezzi et al. (2021) and were obtained by the CKS team (Petigura et al., 2017) with the High Resolution Echelle Spectrometer (HIRES; Vogt et al., 1994) spectrograph attached to the Keck I 10 m telescope (Maunakea, Hawaii). The spectra have high resolution (R 60,000), almost complete wavelength coverage between 3640 Å and 7990 Å, and are publicly available111https://california-planet-search.github.io/cks-website/. Their signal-to-noise ratios (S/Ns) vary from 20 to 300 and have a median value of 60, as measured by Ghezzi et al. (2021).
@@ II.2 Line List
We compiled an initial list of 231 lines for the 16 species analyzed in this work (Na I, Mg I, Al I, Si I, Ca I, Sc I, Sc II, Ti I, Ti II, V, Cr I, Cr II, Mn I, Co I, Ni I, Cu I) from the following sources: Meléndez et al. (2014), Bedell et al. (2014), Schuler et al. (2015), Teske et al. (2015), Bedell et al. (2018) and Liu et al. (2018). Their atomic parameters were retrieved from the Vienna Atomic Line Database (Piskunov et al., 1995; Ryabchikova et al., 1997; Kupka et al., 1999, 2000; Ryabchikova et al., 2015; Pakhomov et al., 2019). For Sc, V, Mn, Co, and Cu, we considered the hyperfine structure and isotopic splittings. We visually inspected all lines in the Solar Atlas spectrum from Hinkle et al. (2000), which is the one originally published by Kurucz et al. (1984) but corrected for telluric absorption (Meléndez et al., 2006), and removed eight lines that were too weak, blended, or close to strong lines (e.g., H).
We automatically measured the equivalent widths (EWs) for the remaining lines using ARES v2 (Sousa et al., 2007, 2015) with the following input parameters: smoothder = 4, space = 3.0, rejt = 0.999, lineresol = 0.1, and miniline = 5. We noted that ARES did not provide good fits for some lines, so we manually measured their EWs using the splot task in the Community Distribution of IRAF (Tody, 1986, 1993; National Optical Astronomy Observatories, 1999; Observatories and community, 2025)222Image Reduction and Analysis Facility (IRAF) was written by the National Optical Astronomy Observatories (NOAO) and is now maintained by the iraf-community. IRAF is listed in the Astronomical Source Code Library as ascl:9911.002. The DOI is 10.5281/zenodo.5816743..
Using this revised line list and the 2019 version of the software MOOG (Sneden, 1973)333https://www.as.utexas.edu/~chris/moog.html, we determined the chemical abundances for the Sun. The model atmosphere was adopted from the KURUCZ ATLAS9 ODFNEW grid (Castelli and Kurucz, 2003) and has the following parameters: = 5777 K, = 4.44, [Fe/H] = 0.00 and = 1.00 km s-1. We used the driver blends on MOOG for Sc, V, Mn, Co, and Cu and abfind for the other elements, both with option 1 for the treatment of damping, i.e., van der Waals damping factors within the code (Barklem et al., 2000; Barklem and Aspelund-Johansson, 2005) are adopted if available; otherwise, values from our line list are utilized. After checking the results, we adjusted the gf values so that all lines for a given element returned the corresponding solar abundances from Asplund et al. (2009). Finally, we removed 18 lines that are located in the interorder or inter-CCD spaces in the HIRES spectra. Our final list contains 205 lines and can be seen in Table 2, which also contains the measured EWs for the Solar Atlas.
@@ II.3 Chemical Abundances
The chemical abundances for the stars in our sample were obtained in a similar manner as for the Solar Atlas. The model atmospheres were interpolated from the KURUCZ ATLAS9 ODFNEW grid (Castelli and Kurucz, 2003) using the atmospheric parameters (, , [Fe/H], and ) from Martinez et al. (2019) and Ghezzi et al. (2021). The EWs were automatically measured with ARES v2 (Sousa et al., 2007, 2015) with the same input parameters described in Section II.2, except for the rejt parameter, which was replaced by the S/N value of the spectrum of each star.
@@
Note, however, that EWs could not be measured for all lines because some of them were too weak or blended, depending on the stellar parameters and S/N of the spectra. In particular, ARES was not able to measure any EWs for Sc I for seven stars: KOI-244, KOI-262, KOI-1530, KOI-2833, KOI-3165, KOI-3438, and KOI-3928. They are hotter stars with lower metallicities, so Sc I lines are very weak (EW 5 mÅ).
@@
For each star, we removed lines with measured depths lower than zero or higher than one, as these are nonphysical values. Additionally, we cut lines with relative errors on the EWs larger than 50%, which indicate poor fits to the profiles, or with EW 150 mÅ, to avoid very strong lines. The cut based on the relative errors left no Al lines for the stars KOI-49, KOI-620, KOI-1779, and KOI-2022. All EWs used for the following abundance determination are provided in Table 2.
@@
We determined the chemical abundances using the driver blends of the 2019 version of MOOG (Sneden, 1973) for elements that have hyperfine structure and isotopic splittings (Sc, V, Mn, Co, and Cu) and abfind otherwise, both with option 1 for the damping input parameter. For species that had two or more lines, we made an initial cut of lines having abundances with absolute deviations from the median larger than 0.40 dex to remove lines with possible problems in their EW measurements. This procedure left no Na lines for the star KOI-837. Then, we performed a cut of lines for which abundances were outside a two-median absolute deviation (MAD) interval around the median. However, we noted that this procedure was not enough to remove bad measurements in all cases, so we added two additional clippings to remove abundances with absolute deviations greater than 0.20 dex. This iterative process ensured that we ended up with precise abundances even for stars with the lowest S/N spectra. Finally, we subtracted the mean abundances from the reference solar values given by Asplund et al. (2009) to obtain [X/H] values, where X is a given element. The abundances for all stars in our sample are shown in Table 1, which also includes the values obtained from the analysis of the spectrum of sunlight reflected off Vesta (see Section III.2.2).
@@
We calculated the abundance uncertainties considering the contributions from the dispersions in the abundances returned by the individual lines as well as the variations caused by the uncertainties in each of the four atmospheric parameters (, , [Fe/H], and ). The former was estimated from the standard deviation of the mean for the absolute abundance A(X), if the element X had abundance measurements from two or more lines. The contribution of the effective temperature for the uncertainty was obtained by generating two new model atmospheres, with (), and determining two new sets of abundances for all elements. Then, we calculated the differences A(X)final - A(X), and the maximum absolute value was taken as the uncertainty caused by the error on . A similar procedure was done for estimating the uncertainties caused by the errors in , [Fe/H], and . The total uncertainties for the abundances of each element were determined by adding all five contributions in quadrature and are shown in Table 1.